Here are simplified financial statements for Phone Corporation in a recent year:INCOME STATEMENT(Figures in $ millions

Net sales$12,600

Cost of goods sold3,700

Other expenses4,127

Depreciation2,338

Earnings before interest and taxes (EBIT)$2,375

Interest expense 655

Income before tax$1,720

Taxes (at 35%)602

Net income$1,118

Dividends$816

BALANCE SHEET(Figures in $ millions)End of YearStart ofYearAssets

End of year start of the year

Cash and marketable securities $83 $152

Receivables 2,082 2,370

Inventories 157 208

Other current assets 837 902

Total current assets 3,159 3,632

Net property, plant, and equipment 19,913 19,855

Other long-term assets 4156 3710

Total assets 27228 27197

Liabilities and shareholders’ equity

Payables 2504 2980

Short-term deb 1389 1543  

Other current liabilities 781 757

Total current liabilities 4674 5280

Long-term debt and leases 8512 7907

Other long-term liabilities 6118 6089

Shareholders’ equity 7924 7921

Total liabilities and shareholders’ equity $27,228 $27,197

Calculate the following financial ratios for Phone Corporation: (Use 365 days in a year. Do not round intermediate calculations. Round your final answers to 2 decimal places.)

a.Return on equity (use average balance sheet figures)

b.Return on assets (use average balance sheet figures)

c.Return on capital (use average balance sheet figures)

d.Days in inventory (use start-of-year balance sheet figures)

e.Inventory turnover (use start-of-year balance sheet figures)

f.Average collection period (use start-of-year balance sheet figures)

g.Operating profit margin

h.Long-term debt ratio (use end-of-year balance sheet figures)

i.Total debt ratio (use end-of-year balance sheet figures)

j.Times interest earned3

k.Cash coverage ratio

l.Current ratio (use end-of-year balance sheet figures)

m.Quick ratio (use end-of-year balance sheet figures)
Answer #1

As per rules I am answering the first 4 subparts of the question

1: ROE = Net Income/ Average Equity

= 1118/((7924+7921)/2)

=14.11%

2: ROA = Net Income/ Average Assets

= 1118/((27228+27197)/2)

=4.11%

3: Return on capital = EBIT/ (Total assets – current liabilities)

= 2375/ (27228-4674)

=10.53%

4: Days in inventory = Ending inventory*365/COGS

= 157*365/ 3700

=15.49 days